OpenStudy (anonymous):

Help please! Will give metal! The volume of a cube is 36 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer. I know the volume of a cube is s^3 and the volume of a pyramid is V=1/3(s^2)(h) but how do I use them to solve this problem.

OpenStudy (anonymous):

@FibonacciChick666 @inkyvoyd @dan815 can any of you help?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Vp=1/3.A.h(i) Vc=L³->L=(36)^(1/3) Vp=1/3.L².L->Vp=36/3=12

OpenStudy (anonymous):

you're a life saver thanks a bunch